# TSF-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting

### v2

@@ Abstract
Time series forecasting aims to model temporal dependencies among variables for future state inference, holding significant importance and widespread applications in real-world scenarios. Although deep learning-based methods have achieved remarkable progress, they still exhibit suboptimal performance in long-term forecasting and data-scarce scenarios. Recent research demonstrates that large language models (LLMs) achieve promising performance in time series forecasting. However, we find existing LLM-based methods still have shortcomings: (1) the absence of a unified paradigm for textual prompt formulation and (2) the neglect of modality discrepancies between textual prompts and time series. To address this, we propose TSF-Prompt, an LLM-based time series forecasting framework integrating multi-prompt information and cross-modal semantic alignment. Specifically, we first construct a unified textual prompt paradigm containing learnable soft prompts and textualized hard prompts. Second, to enhance LLMs’ comprehensive understanding of the forecasting task, we design a semantic space embedding and cross-modal alignment module to achieve cross-modal fusion of temporal and textual information. Finally, the transformed time series from the LLMs are projected to obtain the forecasts. Furthermore, we focus on applying time series forecasting to carbon emissions, aiming to provide a modest contribution to global carbon neutrality. Comprehensive evaluations on 6 public datasets and 3 carbon emission datasets demonstrate that TSF-Prompt is a powerful framework for time series forecasting.
Codehttps://github.com/SanMuGuo/Time-Prompt
@@ Introduction
Time series forecasting, which utilizes past observations to predict future data, finds wide applications in domains such as electricity, weather, and carbon emissions (Wang et al. 2024b). Over the past five years, neural networks with diverse architectures—including linear models (Zeng et al. 2023; Wang et al. 2024a), convolutional neural networks (Wang et al. 2023; Wu et al. 2023), recurrent neural networks (Gu and Dao 2023), and Transformers (Vaswani et al. 2017)—have been applied to time series forecasting. However, existing models face two critical challenges: On one hand, they rely heavily on large-scale labeled samples for parameter optimization, often suffering from limited generalization capability in cross-domain scenarios, which results in performance barriers under few-shot or zero-shot forecasting settings (Gruver et al. 2023; Liu et al. 2024b). On the other hand, existing methods struggle to capture long-term temporal dependencies, resulting in suboptimal long-term forecasting performance (Chen et al. 2024; Ni et al. 2025).
@@
Pretrained foundation models, such as Large Language Models (LLMs), have demonstrated strong few-shot and zero-shot learning capabilities in computer vision and natural language processing. With the technological advancements driven by LLMs, increasing research efforts focus on leveraging their potential for time series forecasting (Shyalika et al. 2024; Jin et al. 2023b; Zhang et al. 2024). However, due to the modality gap between LLMs and time series data, knowledge transfer is required. As shown in Figure 1, the most common current approaches are: a) Time series-based LLMs: achieving modality alignment from time series to the semantic space of LLMs via learnable embedding layers (Zhou et al. 2023); b) Prompt-based LLMs: converting temporal data into textual modality through prompt engineering (Liu et al. 2024b; Xue and Salim 2023); c) Multimodal-based LLMs: using both time series and textual information as multimodal inputs, i.e. a hybrid architecture of the above two strategies (Cao et al. 2023; Jia et al. 2024).
@@
Figure 1: (a) Time series-based LLMs: Lack textual information, failing to leverage LLMs’ potential for time series forecasting. (b) Prompt-based LLMs: Struggle to capture fine-grained temporal patterns from textual prompts. (c) Multimodal-based LLMs: Use both text and time series as inputs but lack cross-modal alignment. (d) Our Method: Integrate multi-prompt information with time series through cross-modal alignment.
Current state-of-the-art approaches predominantly employ multimodal-based LLMs, yet notable limitations persist. On one hand, existing prompt engineering strategies utilize either hard prompts or soft prompts exclusively, failing to explore their synergy for further enhancing LLM performance (Jin et al. 2023a; Pan et al. 2024). For hard prompts, they either solely utilize statistical information of time series or directly textualize raw time series (Liu et al. 2024b). Overall, a unified prompt paradigm remains absent. On the other hand, most works employ simple concatenation for multimodal data fusion, inadequately addressing the semantic gap and feature redundancy across different modalities (Liu et al. 2024a, 2025). Notably, these technical limitations directly lead to significant shortcomings in long-term dependency modeling and noise robustness.
To address the aforementioned technical challenges, inspired by recent research (Liu et al. 2024a; Cao et al. 2023; Jin et al. 2023a), this paper proposes an LLM-empowered time series forecasting framework based on a collaborative mechanism of hard and soft prompts, termed TSF-Prompt. The core innovations of this framework lie in two aspects: a) constructing a unified paradigm for soft-hard prompt synergy and b) establishing a cross-modal information fusion network architecture. Specifically, we first design a dual-path prompt template integrating hard prompts (including temporal statistical features and raw numerical values) and adaptive soft prompts (optimized via gradient backpropagation) based on existing time series data. Then, we build a cross-modal data fusion module, including a semantic space embedding module and a cross-modal alignment module, to embed raw time series data and dual-path prompt texts into the LLMs’ dimension followed by modality adaptation. Finally, the LLMs’ outputs are projected to obtain time series forecasting results. Experimental results demonstrate that TSF-Prompt is effective for time series forecasting and verifies LLMs as capable few-shot and zero-shot time series learners.
Furthermore, given that carbon emissions have become a globally critical issue, we conduct a cross-scale study on carbon emission forecasting—spanning global, national, and urban levels—to provide technical support for the strategic requirements of global carbon neutrality goals. The main contributions of this study are summarized as follows:
  • We propose the TSF-Prompt framework, incorporating a soft-hard prompt synergy mechanism and cross-modal data fusion modules to address limitations in long-term dependency modeling and cross-domain generalization.
  • We develop a synergistic mechanism integrating hard prompts and soft prompts, advancing the establishment of a unified prompt engineering paradigm for time series.
@@
  • We construct a semantic space embedding module and a cross-modal alignment module to enable cross-modal fusion of temporal data and textual information, enhancing feature representation.
@@
  • Extensive experiments on 6 benchmark datasets and 3 carbon emission datasets demonstrate that TSF-Prompt achieves superior performance, particularly in few-shot and zero-shot scenarios.
@@
Figure 2: TSF-Prompt framework. First, hard prompts and soft prompts are constructed based on time series. Second, the time series and prompts are embedded into the LLM’s semantic space, followed by cross-modal alignment. Finally, the fused information is fed into the pretrained LLM and projected through output layers to generate forecasting results.
@@
Recently, leveraging the cross-domain transfer capabilities of LLMs, they have demonstrated exceptional performance in time series forecasting tasks, significantly outperforming traditional methods, especially in few-shot scenarios (Liu et al. 2024c). As shown in Figure 1, the most prevalent current approaches include: a) Time series-based LLMs; b) Prompt-based LLMs; c) Multimodal-based LLMs. Time series-based LLMs achieve cross-modal alignment from time series to the LLM semantic space via learnable embedding layers, encoding temporal data into LLM-readable representations. Prompt-based LLMs convert numerical time series into textual information. Current research focuses on Multimodal-based LLMs, which perform cross-modal semantic alignment between temporal and textual data to overcome inherent limitations of single-modality representations. However, challenges remain in activating LLMs’ forecasting capabilities. Specifically, prompt engineering lacks a unified paradigm: TEMPO and S2IP-LLM solely employ learnable soft prompts, neglecting the importance of hard prompts (Cao et al. 2023; Pan et al. 2024); LLM-Time directly textualizes temporal data as hard prompts for LLMs, resulting in feature redundancy due to high overlap with raw temporal data (Gruver et al. 2023); Time-LLM is limited to constructing hard prompts using only statistical features of temporal data (Jin et al. 2023a). Furthermore, existing multimodal alignment methods rely on feature concatenation, failing to effectively bridge the modality gap (Kowsher et al. 2024; Hu et al. 2025).
@@
As illustrated in Figure 2, TSF-Prompt comprises four components: dual-path prompting, semantic space embedding, cross-modal alignment, and time series forecasting. First, hard and soft prompts are constructed based on temporal data. Subsequently, semantic space embedding is performed through the pre-trained LLM and patch reprogramming module. Since soft prompts are inherently generated within this space, no additional embedding is required. Next, cross-modal alignment is executed. Finally, the fused information is fed into the pre-trained LLM and projected through output layers to generate forecasting results.
@@
Soft Prompts. Soft prompts include prompt pool construction, similarity computation, and Top-k prompt selection. The core idea is to dynamically select soft prompt vectors most relevant to the time series (Cao et al. 2023; Pan et al. 2024). First, initialize a prompt pool. Then, compute the similarity between the time series data and prompt keys in the pool, selecting the Top-k most relevant prompts. Notably, the prompt pool is adaptively updated during iterations.
@@
Hard Prompts. Hard prompts provide a direct yet effective approach to activate forecasting capabilities in LLMs. However, a well-considered unified paradigm remains absent. Based on our observation, existing works either solely textualize temporal data or construct text using only statistical features of time series, failing to integrate both approaches (Jin et al. 2023a; Liu et al. 2024a). Experimental practice indicates this limits performance gains. To address this, we propose a unified hard prompt paradigm. As illustrated in Figure 3, our paradigm comprises four key components: (1) Dataset Description, (2) Task Description, (3) Historical Data, and (4) Statistical Features. Dataset description provides background context for LLMs, while task description, historical data, and statistical features collectively enrich temporal information to enhance the LLMs’ forecasting capabilities. Subsequent experiments demonstrate that both auxiliary information [(1), (2), (3)] and temporal data [(4)] are all effective.
@@
ETTh1 0.433 0.438 $0.449$ 0.445 $0.449$ $0.447$ $0.520$ $0.492$ $0.460$ $0.455$ $0.460$ $0.457$ $0.477$ $0.475$ $0.549$ $0.530$ $0.464$ $0.455$ 0.446 $0.447$
ETTh2 0.369 0.398 0.380 0.403 $0.399$ $0.416$ $0.435$ $0.447$ $0.408$ $0.421$ $0.564$ $0.519$ $0.446$ $0.457$ $1.920$ $1.073$ $0.383$ $0.406$ $0.382$ $0.411$
ETTm1 0.381 0.399 $0.421$ $0.413$ $0.400$ $0.410$ $0.466$ $0.450$ $0.408$ $0.416$ $0.404$ $0.408$ $0.53$ $0.496$ $0.697$ $0.605$ $0.407$ $0.412$ 0.390 0.405
ETTm2 0.283 0.329 $0.291$ $0.335$ $0.294$ $0.336$ $0.331$ $0.365$ $0.299$ 0.333 $0.354$ $0.401$ $0.326$ $0.364$ $1.725$ $0.813$ $0.291$ $0.334$ 0.293 $0.334$
Exchange 0.355 0.400 $0.361$ 0.403 $0.448$ $0.451$ $0.761$ $0.584$ $0.425$ $0.447$ 0.340 $0.414$ $0.843$ $0.596$ $0.757$ $0.645$ $0.364$ $0.405$ $0.391$ $0.417$
Weather 0.254 0.279 $0.266$ $0.285$ $0.257$ $0.284$ $0.288$ $0.314$ $0.258$ $0.285$ $0.265$ $0.317$ $0.337$ $0.379$ $0.259$ $0.320$ $0.261$ $0.281$ 0.255 0.278
munichCarbon $0.412$ $0.465$ $0.411$ $0.465$ $0.433$ $0.486$ $0.485$ $0.518$ 0.415 0.469 $0.394$ $0.456$ $0.611$ $0.599$ 0.358 0.437 $0.412$ $0.465$ $0.430$ $0.477$
chinaCarbon 0.584 0.540 $0.627$ $0.557$ $0.659$ $0.586$ $0.660$ $0.592$ $0.665$ $0.588$ $0.689$ $0.642$ $0.642$ $0.608$ $1.177$ $0.834$ $0.667$ $0.579$ 0.604 0.546
worldCarbon $0.646$ $0.563$ $0.649$ $0.565$ 0.546 0.518 $0.687$ $0.614$ $0.751$ $0.639$ $0.754$ $0.698$ $0.653$ $0.603$ $0.619$ 0.539 $0.735$ $0.616$ 0.615 $0.547$
@@
ETTh1 0.461 0.452 0.477 $0.460$ $0.549$ $0.504$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ $0.510$ $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.477$ 0.455
ETTh2 0.383 0.404 $0.396$ $0.414$ $0.416$ $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ 0.391 0.412
ETTm1 0.411 0.408 $0.424$ $0.411$ $0.425$ $0.427$ $0.511$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ 0.414 0.406
ETTm2 0.288 0.332 $0.297$ $0.339$ $0.302$ $0.342$ $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ 0.290 0.333
Exchange 0.339 0.397 $0.368$ 0.407 $0.423$ $0.448$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.136$ $0.411$ $0.442$ $0.373$ $0.411$
Weather $0.270$ $0.289$ 0.269 0.288 0.255 0.282 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.339$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.273$ $0.289$
munichCarbon 0.412 0.465 $0.426$ $0.477$ $0.434$ $0.486$ $0.513$ $0.538$ $0.445$ $0.491$ $0.413$ $0.475$ $0.605$ $0.609$ 0.402 0.474 $0.443$ $0.491$ $0.423$ $0.474$
chinaCarbon 0.659 0.577 $0.714$ $0.606$ $0.690$ $0.596$ $0.785$ $0.651$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.043$ $0.833$ $0.792$ $0.645$ 0.665 0.580
worldCarbon $0.732$ $0.622$ $0.802$ $0.669$ 0.720 0.607 $0.950$ $0.733$ $0.959$ $0.740$ $0.845$ $0.699$ $0.781$ $0.659$ $0.966$ $0.782$ $0.937$ $0.727$ 0.725 0.609
@@
Semantic Space Embedding. This component primarily involves embedding time series data and hard prompts (Liu et al. 2024a; Hu et al. 2025; Liu et al. 2025). For soft prompts, they are generated directly in the embedding space and thus require no embedding.
Time Series. First, segment the time series into patches, then project them into the LLM’s embedding space to obtain $\hat{X}\in{R^{B\times N\times P\times D}}$, where P is the number of patches and $D$ is the embedding dimension. Given the pre-trained LLM’s vocabulary $E\in{R^{V\times D}}$, where $V$ is the vocabulary size, we realign time series data with the vocabulary through reprogramming, achieving textual representation of temporal data to enhance LLM comprehension. Specifically, learnable projection matrices generate query matrices $Q=\hat{X}{W^{Q}}$, key matrices $K=E{W^{K}}$, and value matrices $V=E{W^{V}}$ for each attention head. Then, multi-head self-attention produces the reprogramming temporal representations.
$\tilde{X}={\rm{MHSA}}(Q,K,V)\in{R^{B\times N\times P\times D}}$ (3)
Hard Prompt. Given the hard prompt $T$, map it to discrete tokens using the LLM’s pretrained tokenizer. Then, convert these discrete tokens into continuous vector representations via the LLM’s input embedding layer. During this process, both the tokenizer and LLM remain frozen with no parameter updates during training.

### v3

# Time-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting

@@ Abstract
Time series forecasting aims to model temporal dependencies among variables for future state inference, holding significant importance and widespread applications in real-world scenarios. Although deep learning-based methods have achieved remarkable progress, they still exhibit suboptimal performance in long-term forecasting. Recent research demonstrates that large language models (LLMs) achieve promising performance in time series forecasting, but this progress is still met with skepticism about whether LLMs are truly useful for this task. To address this, we propose Time-Prompt, a framework for activating LLMs for time series forecasting. Specifically, we first construct a unified prompt paradigm with learnable soft prompts to guide the LLM’s behavior and textualized hard prompts to enhance the time series representations. Second, to enhance LLM’ comprehensive understanding of the forecasting task, we design a semantic space embedding and cross-modal alignment module to achieve fusion of temporal and textual data. Finally, we efficiently fine-tune the LLM’s parameters using time series data. Furthermore, we focus on carbon emissions, aiming to provide a modest contribution to global carbon neutrality. Comprehensive evaluations on 6 public datasets and 3 carbon emission datasets demonstrate that Time-Prompt is a powerful framework for time series forecasting.
@@ Introduction
Time series forecasting, which utilizes past observations to predict future data, finds wide applications in domains such as electricity, weather, and carbon emissions (Wang et al. 2024b). Over the past five years, neural networks with diverse architectures—including linear models (Zeng et al. 2023; Wang et al. 2024a), convolutional neural networks (Wang et al. 2023; Wu et al. 2023), recurrent neural networks (Gu and Dao 2023), and Transformers (Vaswani et al. 2017)—have been applied to time series forecasting. However, existing models face two critical challenges: On one hand, they rely heavily on large-scale labeled samples for parameter optimization, often suffering from limited generalization capability in cross-domain scenarios, which results in performance barriers under few-shot or zero-shot forecasting settings (Gruver et al. 2023; Liu et al. 2024a). On the other hand, existing methods struggle to capture long-term temporal dependencies, resulting in suboptimal long-term forecasting performance (Chen et al. 2024; Ni et al. 2025).
@@
Pretrained foundation models, such as Large Language Models (LLMs), have demonstrated strong few-shot and zero-shot learning capabilities in computer vision and natural language processing. As pre-trained models advance, research is increasingly split between harnessing LLMs for time series forecasting (Shyalika et al. 2024; Jin et al. 2023; Zhang et al. 2024) and training time series foundation models (Goswami et al. 2024; Ansari et al. 2024; Wiliński et al. 2024). Considering the high computational cost of foundation models, our work focuses on the former. As shown in Figure 1, the most common LLMs-based approaches are: (a) Time series-based LLMs: Projecting the time series into the LLM’s semantic space via an embedding layer (Zhou et al. 2023); (b) Prompt-based LLMs: converting temporal data into textual modality through prompt engineering (Liu et al. 2024a); (c) Multimodal-based LLMs: Processing of time series and text separately (Zhou et al. 2025); (d) Multimodal-based LLMs: Feeding unified multimodal information into the LLMs (Zhao, Chen, and Sun 2025).
@@
Figure 1: a) Time series-based LLMs: Lack textual information, failing to leverage LLMs’ potential for time series forecasting. (b) Prompt-based LLMs: Struggle to capture fine-grained temporal patterns from textual prompts. (c) Multimodal-based LLMs: Time series are processed via an encoder, and text via LLMs, separately. (d) Multimodal-based LLMs: Time series and text information are processed by LLMs in a unified manner.
Currently, research primarily focuses on two main types (c) and (d). For example, TimeCMA first encodes the multimodal data, where the time series is processed by an encoder and the text by a frozen LLM. Subsequently, cross-modal alignment is performed, and finally, the prediction result is obtained through a decoder (Liu et al. 2025a). In contrast, Time-LLM processes both time series and text uniformly through a frozen LLM, which then projects the target value (Jin et al. 2024).
Although current methods have achieved state-of-the-art performance, there remains some skepticism—do LLM truly benefit time series forecasting? We believe that with appropriate guidance, the time series forecasting potential of LLMs can be activated. This is because, for LLMs, textual reasoning is fundamentally a form of discrete-value reasoning. Viewed from this perspective, it is quite similar to the task of time series forecasting, with the main distinction being that time series consists of continuous values. Furthermore, LLMs have exhibited certain mathematical reasoning abilities, which also indicates a capacity for handling continuous values.
To unlock the potential of LLMs for time series forecasting, inspired by recent research (Liu et al. 2025a; Cao et al. 2024; Pan et al. 2024; Jin et al. 2024), we propose Time-Prompt, an LLM-empowered framework that activates the potential of LLMs for time series forecasting through time series reprogramming, dual-path prompting, cross-modal alignment, and LoRA. We design a dual-path prompting framework that activates the LLM’s specific patterns with soft prompts and enhances its time series representations with hard prompts. We also reprogram the time series to endow it with a text-like representation. Furthermore, we construct a cross-modal alignment module to eliminate the modality gap between text and time, and we efficiently fine-tune the LLM with LoRA to enhance its ability to process continuous values. In addition, given that carbon emissions have become a globally critical issue, we conduct a cross-scale study on carbon emission forecasting—spanning global, national, and urban levels—to provide technical support for the strategic requirements of global carbon neutrality goals. The main contributions of this study are summarized as follows:
  • We propose Time-Prompt, a framework that activates the potential of LLMs in time series forecasting through strategies such as dual-path prompting, cross-modal alignment, and efficient fine-tuning.
  • We design a dual-path prompting strategy, with learnable soft prompts guiding the LLM’s behavior and hard prompts extracting enhanced time series representations.
@@
  • We construct a semantic space embedding module and a cross-modal alignment module to enable cross-modal fusion of temporal data and textual information, eliminating the modality gap.
@@
  • Extensive experiments on 6 benchmark datasets and 3 carbon emission datasets demonstrate that Time-Prompt achieves superior performance.
@@
Figure 2: Time-Prompt framework. First, hard prompts and soft prompts are constructed based on time series. Second, the time series and prompts are embedded into the LLM’s semantic space, followed by cross-modal alignment. Finally, the fused information is fed into the pretrained LLM and projected through output layers to generate forecasting results.
@@
Recently, leveraging the cross-domain transfer capabilities of LLMs, they have demonstrated exceptional performance in time series forecasting tasks, significantly outperforming traditional methods, especially in few-shot scenarios (Liu et al. 2024b). As shown in Figure 1, the most prevalent current approaches include: (a) Time series-based LLMs; (b) Prompt-based LLMs; (c) Multimodal-based LLMs; (d) Multimodal-based LLMs. Time series–based LLMs project time series into the semantic space of LLMs via an embedding layer, and subsequently map the embedding vectors to obtain prediction values (Zhou et al. 2023). In contrast, Prompt-based LLMs construct the time series as a text modality (Xue and Salim 2023). Lstprompt is a typical Prompt-based method that proposes a prompt construction technique, guiding the LLM to make predictions via a Chain-of-Thought approach (Liu et al. 2024a).
However, single-modality time series forecasting often suffers from limited representational capacity, leading to suboptimal prediction performance. Therefore, many studies attempt to integrate Type (a) and Type (b) approaches, leveraging multimodal information to overcome the inherent limitations of single-modality methods. Early efforts predominantly fall into Type (d), which uniformly feeds time series and textual inputs into LLMs to obtain improved feature representations. GPT4MTS, a foundational work in this category, embeds time series and simple text prompts into the semantic space, from which the large model derives a joint representation (Jia et al. 2024). Many subsequent works build upon this framework. Time-LLM enhances model comprehension of time series by reprogramming the input time series (Jin et al. 2024). Meanwhile, S²IP-LLM and TEMPO employ learnable soft prompts to mark specific temporal patterns, thereby more effectively activating the large model (Cao et al. 2024; Pan et al. 2024).
Subsequently, with growing skepticism about whether LLMs are truly useful for time series (Tan et al. 2024), some studies shift away from feeding time series into LLMs for representation and instead adopt dedicated encoders—this approach corresponds to Type (c). TimeCMA encodes the time series with an encoder and the text with an LLM, followed by cross-modal alignment, and finally generates the prediction via a unified decoder (Liu et al. 2025a). In a related approach, another study further enhances the textual information using reinforcement learning (Su et al. 2025).
The problem with Type (c) is that the LLM only plays an auxiliary role, contributing little to time series forecasting. The issue with Type (d), however, is the LLM’s inability to handle continuous values like those in time series. Nevertheless, recent research continues to emerge for both Type (c) (Zhou et al. 2025; Su et al. 2025) and Type (d) (Chen et al. 2025; Zhang et al. 2025), in an effort to address the widespread skepticism. Our method belongs to the Type (d). LLMs possess powerful text-based reasoning capabilities. Fundamentally, their primary distinction from time series lies in the nature of data: discrete versus continuous values. Therefore, time series forecasting and textual reasoning exhibit significant similarity. Moreover, LLMs have also demonstrated strong reasoning abilities in mathematics. So we believe that with appropriate guidance, LLMs can enhance the performance of time series forecasting. Our method mainly contributes in this area. On the LLM side, we guide the LLM’s latent patterns with soft prompts and achieve efficient fine-tuning through LoRA. For the time series aspect, we reprogram the time series data and incorporate cross-modal alignment.
@@
The overall framework of Time-Prompt is illustrated in Figure 2. First, we construct a dual-path prompting framework for the time series forecasting task. This framework incorporates both hard prompts (including temporal statistical features and historical data) and learnable soft prompts. Our goal is to guide the LLM’s internal behavioral patterns via soft prompts and enhance time series feature representation through hard prompts. Second, we embed both the time series and text into the LLM’s semantic space via semantic space embedding module to facilitate better comprehension by the LLM. Subsequently, to address the modality gap between text and time series, we employ cross-modal alignment. Finally, the global and local features of the time series are extracted and projected to obtain the predicted values. Here, the LLM is fine-tuned using LoRA to adapt it from discrete-value reasoning to continuous-value prediction.
@@
Soft Prompts. Soft prompts include prompt pool construction, similarity computation, and Top-k prompt selection. The core idea is to dynamically select soft prompt vectors most relevant to the time series. First, initialize a prompt pool. Then, compute the similarity between the time series data and prompt keys in the pool, selecting the Top-k most relevant prompts. Notably, the prompt pool is adaptively updated during iterations.
@@
Hard Prompts. Hard prompts provide a direct and effective way to activate the LLM’s time series forecasting capability, as no modality gap exists between the LLM and the prompts. Hard prompts can also be viewed as enhanced representations of the time series. As illustrated in Figure 3, our paradigm comprises three key components: (1) Instruction, (2) Statistical Characteristics, and (3) Historical Data. The Instruction provides the LLM with the task context, the Statistical characteristics correctly guide the LLM’s patterns, and the Historical Data is the direct textualization of the time series data.
@@
ETTh1 0.430 0.439 $0.449$ 0.445 $0.449$ $0.447$ $0.520$ $0.492$ $0.460$ $0.455$ $0.460$ $0.457$ $0.477$ $0.475$ $0.549$ $0.530$ $0.464$ $0.455$ 0.446 $0.447$
ETTh2 0.368 0.397 0.380 0.403 $0.399$ $0.416$ $0.435$ $0.447$ $0.408$ $0.421$ $0.564$ $0.519$ $0.446$ $0.457$ $1.920$ $1.073$ $0.383$ $0.406$ $0.382$ $0.411$
ETTm1 0.380 0.398 $0.421$ $0.413$ $0.400$ $0.410$ $0.466$ $0.450$ $0.408$ $0.416$ $0.404$ $0.408$ $0.53$ $0.496$ $0.697$ $0.605$ $0.407$ $0.412$ 0.390 0.405
ETTm2 0.281 0.326 $0.291$ $0.335$ $0.294$ $0.336$ $0.331$ $0.365$ $0.299$ 0.333 $0.354$ $0.401$ $0.326$ $0.364$ $1.725$ $0.813$ $0.291$ $0.334$ 0.293 $0.334$
Exchange 0.331 0.390 $0.361$ 0.403 $0.448$ $0.451$ $0.761$ $0.584$ $0.425$ $0.447$ 0.340 $0.414$ $0.843$ $0.596$ $0.757$ $0.645$ $0.364$ $0.405$ $0.391$ $0.417$
Weather 0.252 0.277 $0.266$ $0.285$ $0.257$ $0.284$ $0.288$ $0.314$ $0.258$ $0.285$ $0.265$ $0.317$ $0.337$ $0.379$ $0.259$ $0.320$ $0.261$ $0.281$ 0.255 0.278
munichCarbon 0.411 0.464 $0.411$ $0.465$ $0.433$ $0.486$ $0.485$ $0.518$ $0.415$ $0.469$ $0.394$ $0.456$ $0.611$ $0.599$ 0.358 0.437 $0.412$ $0.465$ $0.430$ $0.477$
chinaCarbon 0.571 0.532 $0.627$ $0.557$ $0.659$ $0.586$ $0.660$ $0.592$ $0.665$ $0.588$ $0.689$ $0.642$ $0.642$ $0.608$ $1.177$ $0.834$ $0.667$ $0.579$ 0.604 0.546
worldCarbon $0.617$ $0.544$ $0.649$ $0.565$ 0.546 0.518 $0.687$ $0.614$ $0.751$ $0.639$ $0.754$ $0.698$ $0.653$ $0.603$ $0.619$ 0.539 $0.735$ $0.616$ 0.615 $0.547$
@@
ETTh1 0.452 0.445 0.477 $0.460$ $0.549$ $0.504$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ $0.510$ $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.477$ 0.455
ETTh2 0.381 0.404 $0.396$ $0.414$ $0.416$ $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ 0.391 0.412
ETTm1 0.394 0.402 $0.424$ $0.411$ $0.425$ $0.427$ $0.511$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ 0.414 0.406
ETTm2 0.284 0.327 $0.297$ $0.339$ $0.302$ $0.342$ $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ 0.290 0.333
Exchange 0.337 0.395 $0.368$ 0.407 $0.423$ $0.448$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.136$ $0.411$ $0.442$ $0.373$ $0.411$
Weather 0.265 0.286 $0.269$ $0.288$ 0.255 0.282 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.339$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.273$ $0.289$
munichCarbon 0.406 0.459 $0.426$ $0.477$ $0.434$ $0.486$ $0.513$ $0.538$ $0.445$ $0.491$ $0.413$ $0.475$ $0.605$ $0.609$ 0.402 0.474 $0.443$ $0.491$ $0.423$ $0.474$
chinaCarbon 0.653 0.575 $0.714$ $0.606$ $0.690$ $0.596$ $0.785$ $0.651$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.043$ $0.833$ $0.792$ $0.645$ 0.665 0.580
worldCarbon 0.705 0.601 $0.802$ $0.669$ 0.720 0.607 $0.950$ $0.733$ $0.959$ $0.740$ $0.845$ $0.699$ $0.781$ $0.659$ $0.966$ $0.782$ $0.937$ $0.727$ $0.725$ $0.609$
@@
Semantic Space Embedding. This component primarily involves embedding the time series and prompts into the LLM’s semantic space to enhance the LLM’s understanding of the variables. For soft prompts, they are generated directly in the embedding space and thus require no embedding.
Time Series. First, segment the time series into patches, then project them into the LLM’s embedding space to obtain $\hat{X}\in{R^{B\times N\times P\times D}}$, where P is the number of patches and $D$ is the embedding dimension. Given the pre-trained LLM’s vocabulary $E\in{R^{V\times D}}$, where $V$ is the vocabulary size, we realign time series data with the vocabulary through reprogramming, achieving textual representation of temporal data to enhance LLM comprehension. Specifically, learnable projection matrices generate query matrices $Q=\hat{X}{W^{Q}}$, key matrices $K=E{W^{K}}$, and value matrices $V=E{W^{V}}$ for each attention head. Then, multi-head cross-attention produces the reprogramming temporal representations.
$\tilde{X}={\rm{MHCA}}(Q,K,V)\in{R^{B\times N\times P\times D}}$ (3)
Hard Prompt. Given the hard prompt $T$, map it to discrete tokens using the LLM’s pretrained tokenizer. Then, convert these discrete tokens into continuous vector representations via the LLM’s input embedding layer. During this process, the LLM’s tokenizer remain frozen with no parameter updates during training.